OpenStudy (anonymous):

95 is 20% of what number

OpenStudy (anonymous):

20X/100=95 x/5=95... where x is the number ur seeking.

OpenStudy (anonymous):

475 it is....

Parth (parthkohli):

Let's say that 95 is 20% of x. \(\Large \color{Black}{\Rightarrow 95 = 0.2x }\) \(\Large \color{Black}{\Rightarrow x = {95 \over 0.2} }\) \(\Large \color{Black}{\Rightarrow x = 475 }\)

OpenStudy (anonymous):

okay :(

Parth (parthkohli):

Do you get it?

OpenStudy (anonymous):

yes i do thank you(:

OpenStudy (anonymous):

there is a lot more just wait... lmao(:

OpenStudy (anonymous):

\[\left(\begin{matrix}95 \\ x \end{matrix}\right)=\left(\begin{matrix}20 \\ 100\end{matrix}\right)\] cross multiply the proportion you get 9500= 20x divide each side by 20 that's your answer

OpenStudy (anonymous):

get it